Arnold and Juju are 140 m apart. Arnold can run at a speed of 8 m/s while Juju can run at a speed of 6m/s. If the two boys run at the same instant, straight towards each other, when and where will they meet?


1
Expert's answer
2020-12-31T08:07:09-0500

They approach each other with a speed "v=v_{Arnold}+v_{Juju} = 8+6=14 m\/s" . Therefore, time when they meet is "t = \frac{distance}{velocity} = \frac{140}{14} = 10s" , 10 seconds. During this time, Arnold will run "10*8 = 80m" from his initial position, so that's where they will meet (we can equivalently say that they will meet 60m from Juju's initial position).